. A mild steel tank of wall thickness 12 mm contains water at 95°C. The

thermal conductivity of mild steel is 50 W/m°C, and the heat transfer coefficients for the inside

and outside the tank are 2850 and 10 W/m2°C, respectively. If the atmospheric temperature is

15°C, calculate :

(i) The rate of heat loss per m2 of the tank surface area ;

(ii) The temperature of the outside surface of the tank


1
Expert's answer
2021-09-04T15:19:44-0400


Thickness of mild steel tank wall

"L=12 \;mm = 0.012 \;m"

Temperature of water

"t_{hf}= 95 \;°C"

Temperature of air

"t_{cf}= 15 \;°C"

Thermal conductivity of mild steel

"k=50 \;W/m°C"

Heat transfer coefficients :

Hot fluid (water)

"h_{hf}= 2850 \;W/m^2°C"

Cold fluid (air)

"h_{cf}= 10 \;W/m^2°C"

(i) Rate of heat loss per m2 of the tank surface area, q :

Rate of heat loss per m2 of tank surface,

"q=UA(t_{hf}-t_{cf})"

The overall heat transfer coefficient, U is found from the relation

"\frac{1}{U}= \frac{1}{h_{hf}}+ \frac{L}{k} + \frac{1}{h_{cf}}= \frac{1}{2850}+ \frac{0.012}{50} + \frac{1}{10} \\

= 0.0003508 + 0.00024 + 0.1 = 0.1006 \\

U= \frac{1}{0.1006}=9.94 \;W/m^2°C \\

q= 9.94 \times 1 \times (95-15) = 795.2 \;W/m^2"

(ii) Temperature of the outside surface of the tank, t2 :

We know that

"q=h_{cf} \times 1 \times (t_2-t_{cf}) \\

975.2=10(t_2-15) \\

t_2=\frac{795.2}{10}+15 = 94.52 \;°C"